# TinyVLA: Towards Fast, Data-Efficient Vision-Language-Action Models for Robotic Manipulation

### v2

@@ I Introduction
Training multitasking robot imitators to operate in complex and uncertain environments faces considerable challenges due to limited data and the difficulty of learning physical motion [1, 2, 3]. Moreover, traditional robot models struggle to adapt to new scenes and tasks and are easily affected by distractors, lighting conditions, and background changes [4, 5]. Modern methods typically leverage off-the-shelf Large Language Models (LLMs) [6, 7, 8, 9, 10, 11] for scene descriptions to generate object affordance, location, or heatmaps, followed by a predefined motion planner to complete the tasks [12, 13, 14, 15, 16].
@@ II Related Works
Multimodal Models connect vision and language and extend the reasoning ability of LLMs to process with multimodal input. Numerous works have been proposed in this direction [20, 21, 22, 23, 24, 25, 26, 27, 28]. These MLLMs typically have parameters ranging from 7B to 70B, making the inference cost-prohibitive and limiting the accessibility of MLLMs to a wider audience. Recently, a select number of studies [29, 30, 31, 32, 33, 34, 35, 36]have delved into the exploration of efficient multimodal, with a number of parameters less than 3B, from diverse angles. Robot learning [4, 37, 17, 38, 39, 40, 41, 42, 43, 44] is an crucial topic in the robotics. A number of works introduce vision-language models to the domain of embodied control, these representative works include RT-2 [17] and OpenVLA [18]. Our paper proposes a compact version of VLA models that does not require large-scale robot data pretraining and enables fast inference.
@@ III-A Building TinyVLA with Efficient Vision-Language Models
The initial step involves acquiring pre-trained multimodal language models. While existing works typically focus on vision-language models with over three billion parameters, we trained a more compact vision-language model with parameters ranging from 70 million to 1.4 billion. Our model utilizes Pythia [46] as the language model backend. We then followed the training pipeline of LLaVA [23], using their vision-language dataset to train this family of VLMs. For robot data fine-tuning, we retained all modules from our VLM, including the visual backbone and the vision-language alignment module.
@@ III-B Robot Data Finetuning for Manipulation
Furthermore, the model must preserve the intrinsic knowledge of the language models. The trainable parameters constitute only 5.0% of the entire Transformer’s parameters. We posit that this approach enables the pre-trained model to process inputs with maximum linguistic fidelity while retaining flexibility. Empirical evidence (in the Appendix) suggests that either fully fine-tuning weights or freezing Transformer layers degrades performance. After training is completed, we apply re-parameterization techniques to integrate the LoRA module seamlessly into the standard language model, thereby enhancing inference speed.
@@ IV-A2 Real Robot Setup
Tasks. In the single-arm setting, there are five tasks: 1) closing the drawer (CloseDrawer), 2) stacking the pink cube on top of the blue cube (StackCubes), 3) opening the lid of the box (OpenBox), 4) placing a tennis ball into the ball box, and 5) uprighting a tipped-over mug (FlipMug). In the bimanual robot experiment, we set up three tasks that involved cooperation between two arms: 1) transferring bread to a plate (TransferBread), 2) unzipping a bag and placing a tennis ball inside it (PlaceTennisBag), and 3)stacking cubes on a plate (StackCubes). It is worth noting that the action spaces of tasks vary considerably. For instance, flip mug necessitates the robot to perform wide-ranging rotations to insert the gripper into the mug laterally, which is completely different from stack cubes which is pick&place type. The span of different trajectories within the same task varies markedly as well, e.g., the length of $stack~{}cubes$ trajectories ranges from 100 to 300. This provides TinyVLA with more challenges in learning to perform these tasks.
@@ IV-C Generalization to Unseen Instructions
In this work, we investigate the generalization capabilities of TinyVLA-H, which demonstrates the best performance in both real-world scenarios and simulations. Since TinyVLA uses a pre-trained multimodal model as its backbone, we observe similar embodied capabilities driven by the rich world knowledge implicitly stored in these models, even though the fine-tuned version is not trained on question-answering pair data like RT-2 [17]. As demonstrated in Figure 4, we evaluated with a fixed list of instructions (i.e., ”Pick the [object]”), where [object] are randomized objects that have not been seen in the training data. We use obj. as the abbreviation of objects in Figure 4. We test with three objects, a mug, a toy car, and a pink cube.
@@
The second level involves grasping the object. Both objects presented have been part of the training data. We asked the model to ”pick the cube”. Despite the environment and instruction not being part of the training data, TinyVLA successfully picked up the cube. This indicates that TinyVLA effectively maps textual descriptions to physical objects.
To further increase the difficulty of the test, we designed the third level, where the model is instructed to ”pick a toy car” and ”place it into the box”. The toy car is not in the training data. We placed a pink cube beside the toy car to assess whether the model could comprehend the instructions. Additionally, the command ”place into the box” introduces a new skill-object combination, suggesting that even though the object is familiar, its function has been altered. Successfully completing this task indicates that TinyVLA possesses the ability to recognize novel objects and identify new functionalities in familiar ones.
@@ IV-D More Real-World Experiments: Bimanual Robot
We further conducted experiments on the Bimanual UR5 Robot, applying it to three distinct tasks: PlaceBread, StackCube, and PlaceTennisBag. These tasks vary significantly in both duration and required skills, posing challenges for training a multi-task policy model. As demonstrated in Table III, the Diffusion Policy achieved a zero success rate across all three tasks. In contrast, our TinyVLA-H model attained an average success rate of 65%. Notably, the OpenVLA fails in every trial. We suspect this is because OpenVLA is pre-trained on the OpenX dataset, which consists entirely of single-arm robot data, making it ineffective when applied to bimanual robots.

### v4

@@ I Introduction
Training multitasking robot imitators to operate in complex and uncertain environments faces considerable challenges due to limited data and the difficulty of learning physical motion [1, 2, 3, 4]. Moreover, traditional robot models struggle to adapt to new scenes and tasks and are easily affected by distractors, lighting conditions, and background changes [5, 6]. Modern methods typically leverage off-the-shelf Large Language Models (LLMs) [7, 8, 9, 10, 11, 12] for scene descriptions to generate object affordance, location, or heatmaps, followed by a predefined motion planner to complete the tasks [13, 14, 15, 16, 17, 18].
@@ II Related Works
Multimodal Models connect vision and language and extend the reasoning ability of LLMs to process with multimodal input. Numerous works have been proposed in this direction [22, 23, 24, 25, 26, 27, 28, 29, 30]. These MLLMs typically have parameters ranging from 7B to 70B, making the inference cost-prohibitive and limiting the accessibility of MLLMs to a wider audience. Recently, a select number of studies [31, 32, 33, 34, 35, 36, 37]have delved into the exploration of efficient multimodal, with a number of parameters less than 3B, from diverse angles. Robot learning [38, 39, 19, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] is an crucial topic in the robotics. A number of works introduce vision-language models to the domain of embodied control, these representative works include RT-2 [19] and OpenVLA [20]. Our paper proposes a compact version of VLA models that does not require large-scale robot data pretraining and enables fast inference.
@@ III-A Building TinyVLA with Efficient Vision-Language Models
The initial step involves acquiring pre-trained multimodal language models. While existing works typically focus on vision-language models with over three billion parameters, we trained a more compact vision-language model with parameters ranging from 40 million to 1.4 billion. Our model utilizes Pythia [63] as the language model backend. We then followed the training pipeline of LLaVA [25], using their vision-language dataset to train this family of VLMs. For robot data fine-tuning, we retained all modules from our VLM, including the visual backbone and the vision-language alignment module.
@@ III-B Robot Data Finetuning for Manipulation
Furthermore, the model must preserve the intrinsic knowledge of the language models. The trainable parameters constitute only 5.0% of the entire transformer’s parameters. We posit that this approach enables the pre-trained model to process inputs with maximum linguistic fidelity while retaining flexibility. After training is completed, we apply re-parameterization techniques to integrate the LoRA module seamlessly into the standard language model, thereby enhancing inference speed.
@@ IV-A2 Real Robot Setup
Tasks. In the single-arm setting, there are five tasks: 1) closing the drawer (CloseDrawer), 2) stacking the pink cube on top of the blue cube (StackCubes), 3) opening the lid of the box (OpenBox), 4) placing a tennis ball into the ball box, and 5) uprighting a tipped-over mug (FlipMug). In the bimanual robot experiment, we set up three tasks that involved cooperation between two arms: 1) transferring bread to a plate (TransferBread), 2) unzipping the bag and placing a tennis ball inside it (PlaceTennisBag), and 3)stacking cubes on a plate (StackCubes). It is worth noting that the action spaces of tasks vary considerably. For instance, flip mug necessitates the robot to perform wide-ranging rotations to insert the gripper into the mug laterally, which is completely different from stack cubes which is pick&place type. The span of different trajectories within the same task varies markedly as well, e.g., the length of $stack~{}cubes$ trajectories ranges from 100 to 300. This provides TinyVLA with more challenges in learning to perform these tasks.
@@ IV-C Generalization to Unseen Instructions
In this work, we investigate the generalization capabilities of TinyVLA-H, which demonstrates the best performance in both real-world scenarios and simulations. Since TinyVLA uses a pre-trained multimodal model as its backbone, we observe similar embodied capabilities driven by the rich world knowledge implicitly stored in these models, even though the fine-tuned version is not trained on question-answering pair data like RT-2 [19]. As demonstrated in Figure 4, we evaluated with a fixed list of instructions (i.e., “Pick the [object]”), where [object] are randomized objects that have not been seen in the training data. We use obj. as the abbreviation of objects in Figure 4. We test with three objects, a mug, a toy car, and a pink cube.
@@
The second level involves grasping the object. Both objects presented have been part of the training data. We asked the model to “pick the cube”. Despite the environment and instruction not being part of the training data, TinyVLA successfully picked up the cube. This indicates that TinyVLA effectively maps textual descriptions to physical objects.
To further increase the difficulty of the test, we designed the third level, where the model is instructed to “pick a toy car” and “place it into the box”. The toy car is not in the training data. We placed a pink cube beside the toy car to assess whether the model could comprehend the instructions. Additionally, the command “place into the box” introduces a new skill-object combination, suggesting that even though the object is familiar, its function has been altered. Successfully completing this task indicates that TinyVLA possesses the ability to recognize novel objects and identify new functionalities in familiar ones.
@@ IV-D More Real-World Experiments: Bimanual Robot
We further conducted experiments on the Bimanual UR5 Robot, applying it to three distinct tasks: PlaceBread, StackCube, and PlaceTennisBag. These tasks vary significantly in both duration and required skills, posing challenges for training a multi-task policy model. As shown in Table III, while the Diffusion Policy excels in the PlaceTennisBag task, our TinyVLA-H model achieved an average success rate of 44.5%, surpassing the Diffusion Policy’s 38.2%. Notably, the OpenVLA fails in every trial. We suspect this is because OpenVLA is pre-trained on the OpenX dataset, which consists entirely of single-arm robot data, making it ineffective when applied to bimanual robots.
@@ V Conclusion
VI Acknowledgements
We sincerely thank Yanjie Ze for contributions in discussions and paper review. Our also thank Yong Wang for his assistance with the hardware setup.